OpenStudy (anonymous):

an article was sold at16% gain. Had it been sold for rupees 200 more.the gain would have been 20% then the cost price of article is?

Directrix (directrix):

Let x be the merchant's cost of the article. Selling price to customer: x + .16x --------------- sold for rupees 200 more x + .16x + 200 = x + .20x x = 5,000 --------------- @suman2335 Please check my equations and determine if you agree with my thinking about this problem. Thanks.

OpenStudy (anonymous):

how these20 came?

Directrix (directrix):

>>>>.the gain would have been 20% then the cost price of article is. 20% = .20 The cost is x so the gain is .20x

OpenStudy (anonymous):

sorry but how these 5000 came out can u explain me?

Directrix (directrix):

an article was sold at16% gain Let x be the merchant's cost of the article. Selling price to customer: x + .16x

Directrix (directrix):

Had it been sold for rupees 200 more: New selling price to customer: x + .16x + 200

Directrix (directrix):

the gain would have been 20% x + .20x

Directrix (directrix):

the cost price of article is? x + .16x + 200 = x + .20x Solve for x.

OpenStudy (anonymous):

and

Directrix (directrix):

x + .16x + 200 = x + .20x 1.16x + 200 = 1.2x and